OpenStudy (anonymous):

HELP WITH 10 QUESTIONS?!?! HERE IS THE FIRST 1. Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4-x and y = 2x + 3 intersect are the solutions of the equation 4-x = 2x + 3. Part B: Make tables to find the solution to 4-x = 2x + 3. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4-x = 2x + 3 graphically? A system of equations is shown below. 5x - 5y = 10 3x - 2y = 2

OpenStudy (anonymous):

Well not 10, maybe a little less

OpenStudy (phi):

what did you say for part A?

OpenStudy (anonymous):

That is the thing I do not understand what they want me to do... @phi

OpenStudy (anonymous):

@mathstudent55 I got part B. Can you help with a and c?? PLease and thankss.

OpenStudy (phi):

for part a) you have the equations of two lines: y = 4-x and y = 2x + 3 If a point is on the first line AND that same point is on the other line, then the point must be at the intersection (where the lines cross). the y value of this point will be BOTH 4-x and 2x+3. In other words 4-x= 2x+3

OpenStudy (phi):

For part c) graph the two lines y = 4-x and y = 2x + 3 and see where they intersect. That is, pick out the (x,y) pair that is on both lines.